OpenStudy (anonymous):

Find the first six terms of the sequence. \[a_{1}=4, a _{n}=2 \times a _{n-1}\] how do i go about solving this?

OpenStudy (harsimran_hs4):

in the general expression put n =2 a2 = 2 * (a1) = 2*4 = 8 similarly you can find a3 a4 and so on clear?

OpenStudy (anonymous):

umm im unclear on how you got a2 = 2 * (a1) = 2*4 = 8

OpenStudy (anonymous):

could you elaborate?

OpenStudy (harsimran_hs4):

an = 2×an−1(look above in the question it`s given) so put n = 2 in this equation and tell me what you get?

OpenStudy (anonymous):

2 x 4-1?

OpenStudy (harsimran_hs4):

no!!

OpenStudy (harsimran_hs4):

 clear??